# Affine Chord Sobolev Inequalities from Integral Geometry

### v1

@@ Abstract.
Functional analogues of isoperimetric inequalities for chord power integrals are established for $\alpha>0$, together with a complete characterization of the equality cases. For $\alpha\in(0,n)$, these inequalities—referred to as chord Sobolev inequalities— serves as a counterpart to the fractional Sobolev inequality of Almgren and Lieb, which corresponds to the functional isoperimetric inequality for chord power integrals with $\alpha\in(-1,0)$.
Moreover, affine isoperimetric inequalities for chord power integrals are derived, that are significantly stronger than their Euclidean counterparts. Extending the framework to functional and affine settings, functional extensions of chord power integrals and radial mean bodies are introduced. The inclusion relations for classical radial mean bodies due to Gardner and Zhang are extended to both log-concave and $s$-concave function, and the extensions are shown to be sharp.
@@ 1. Introduction
For $\alpha>-1$, the chord power integral of a convex body $K\subset\mathbb{R}^{n}$ (a convex compact set with non-empty interior), is defined by
$I_{\alpha+1}(K)=\int_{{\rm Aff}(n,1)}|K\cap l|^{\alpha+1}dl,$
where ${\rm Aff}(n,1)$ is the affine Grassmannian consistent of $1$-dimensional lines in $\mathbb{R}^{n}$, and $dl$ is the Haar measure on it. The quantity $|K\cap l|$ represents the length of the intersection. The chord power integral has attracted increasing attention in recent years, mainly in connection with Minkowski-type problems (see for example [10, 17, 40, 48, 49]). In this paper, we study inequalities from the viewpoint of chord power integrals, while previous studies on related inequalities approached their problems from different perspectives.
The isoperimetric inequalities for chord power integrals (cf. [46]) state that, given a convex body $K\subset\mathbb{R}^{n}$,
$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\geq\Big(\frac{|K|}{|B^{n}|}\Big)^{\frac{n+\alpha}{n}},$ (1.1)
for $\alpha\in(-1,0)\cup(n,\infty)$, and
$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\leq\Big(\frac{|K|}{|B^{n}|}\Big)^{\frac{n+\alpha}{n}},$ (1.2)
for $\alpha\in(0,n)$. Equality holds in both cases if and only if $K$ is a ball. Throughout, $|K|$ denotes the volume of $K$, that is, the $n$-dimensional Lebesuge measure of $K$, and $B^{n}$ stands for the Euclidean unit ball in $\mathbb{R}^{n}$ with volume $\omega_{n}$.
For $\alpha\in(-1,0)$, the chord power integral $I_{\alpha+1}(K)$ is also referred to as the fractional perimeter, which is closely related to the fractional Sobolev norm (cf. [13, 32, 33]):
$\|f\|_{W^{-\alpha,1}(\mathbb{R}^{n})}=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,$
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a measurable function, and $|\cdot|$ denotes the Euclidean norm in $\mathbb{R}^{n}$. Replacing $f$ with the characteristic function $\operatorname{1}_{K}$ of a convex body $K\subset\mathbb{R}^{n}$, which takes value $1$ on $K$ and 0 elsewhere, yields
$I_{\alpha+1}(K)=\frac{-\alpha(\alpha+1)}{n\omega_{n}}\int_{K}\int_{K^{c}}\frac{1}{|x-y|^{n-\alpha}}dxdy,$ (1.3)
which follows from the Blaschke-Petkantschin formula (cf. [46, Theorem 7.2.7]).
Almgren and Lieb [1] established the Pólya–Szegö inequalities for fractional Sobolev norms, and the equality case was subsequently characterized by Frank and Seiringer [13]. A direct consequence is the fractional Sobolev inequality: for $\alpha\in(-1,0)$ and $f\in W^{-\alpha,1}(\mathbb{R}^{n})$,
$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,$ (1.4)
where $\|f\|_{p}$ denotes the $L^{p}$ norm of $f$ and $W^{-\alpha,1}(\mathbb{R}^{n})$ denotes the fractional Sobolev space of $L^{1}$ functions with finite $\|f\|_{W^{-\alpha,1}(\mathbb{R}^{n})}$. Equality holds in (1.4) if and only if $f$ is a constant multiple of the characteristic function of a ball. The best constant $\sigma_{n,\alpha}$ is given by
$\sigma_{n,\alpha}=\frac{nI_{\alpha+1}(B^{n})}{|\alpha|(\alpha+1)\omega_{n}^{\frac{\alpha}{n}}}=\frac{n2^{\alpha}\pi^{\frac{n-\alpha-1}{2}}\Gamma(\frac{\alpha+1}{2})\Gamma(\frac{n}{2}+1)^{\frac{\alpha}{n}}}{|\alpha|\Gamma(\frac{n+\alpha}{2}+1)},$ (1.5)
where $\Gamma$ denotes the gamma function. The constant can be obtained from the results in [12] as described in [16].
Given the sharp fractional Sobolev inequality introduced above—together with its best constant and equality cases—and the geometric isoperimetric inequality for chord power integrals, it becomes natural to regard (1.4) as a functional analogue of (1.1) for $\alpha\in(-1,0)$. More precisely, replacing $f$ with the characteristic function $\operatorname{1}_{K}$ of a convex body $K\subset\mathbb{R}^{n}$ reduces the inequality to the isoperimetric inequality for $I_{\alpha+1}(K)$ with $\alpha\in(-1,0)$.
Building on this correspondence, it is natural to ask whether the isoperimetric inequalities established for $\alpha\in(0,n)$ and $\alpha>n$ also admit functional counterparts.
Question. Does there exist a functional inequality for a suitable class of functions $f$ whose geometric instance, corresponding to $f=\operatorname{1}_{K}$, recovers the isoperimetric inequalities for chord power integrals with $\alpha>0$?
Noting that for $\alpha>0$, the quantity $I_{\alpha+1}(K)$ also admits a double integral representation, as shown in [40, Lemma 3.3]:
$I_{\alpha+1}(K)=\frac{\alpha(\alpha+1)}{n\omega_{n}}\int_{K}\int_{K}\frac{1}{|x-y|^{n-\alpha}}dxdy.$
This representation suggests a connection with the Hardy-Littlewood-Sobolev inequality (HLS inequality) established by Lieb [30]: for $0<\alpha<n$ and non-negative $f\in L^{p}(\mathbb{R}^{n})$ with $p=2n/(n+\alpha)$,
$\gamma_{n,\alpha}\|f\|_{\frac{2n}{n+\alpha}}^{2}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{f(x)f(y)}{|x-y|^{n-\alpha}}dxdy.$ (1.6)
Equality in (1.6) occurs if and only if $f(x)=a(1+\lambda|x-x_{0}|^{2})^{-(n+\alpha)/2}$ for $x\in\mathbb{R}^{n}$ with $a\geq 0,\lambda>0$ and $x_{0}\in\mathbb{R}^{n}$. This observation leads to the unexpected conclusion that the HLS inequality is not the functional counterpart of (1.2). On further inspection, the essential obstruction appears to be the $2$-homogeneity of the HLS inequality in $f$. Motivated by the 1-homogeneous structure arising in the case $\alpha\in(-1,0)$, we turn to a formulation that reflects the intrinsic homogeneity underlying the geometric setting. This leads naturally to the following 1-homogeneous analogue of (1.6).
Theorem 1 (Chord Sobolev inequalities).
For $\alpha\in(0,n)$ and non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy,$
where the sharp constant $\sigma_{n,\alpha}$ is given by (1.5). Equality holds if and only if $f$ is a constant multiple of the characteristic function of a ball.
With the double integral appearing in Theorem 1 and the fractional Sobolev norm (1.3), we can naturally extend the notion of the chord power integral to functions $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ by setting
$I_{\alpha+1}(f)=\frac{-\alpha(\alpha+1)}{2n\omega_{n}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,\quad\quad\alpha\in(-1,0),$
and
$I_{\alpha+1}(f)=\frac{\alpha(\alpha+1)}{n\omega_{n}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy,\quad\quad\alpha>n.$
A more unified formulation can be written as
$I_{\alpha+1}(f)=\int_{0}^{\infty}I_{\alpha+1}(\{f\geq r\})dr,$
which follows from the coarea formula when $\alpha\in(-1,0)$ and from the layer-cake representation of $\min\{f(x),f(y)\}$ when $\alpha>n$. Here $f$ is typically assumed to be an integrable log-concave function, although the log-concavity assumption is not required in general. For instance, when $\alpha\in(-1,0)$, it suffices that $f$ is a non-negative function in the corresponding fractional Sobolev space.
The inequality in Theorem 1 is a functional counterpart of the classical isoperimetric inequality (1.2). It is invariant under translations and rotations, but not under volume-preserving linear transformations. Affine inequalities, which remain invariant under translations and volume-preserving linear transformations, have proved to be very powerful, as they often strengthen the corresponding Euclidean inequalities. A famous example is the Petty projection inequality for convex bodies in $\mathbb{R}^{n}$. It strengthens the Euclidean isoperimetric inequality and also directly implies it. See, for example, [14, 45], and see [7, 19, 23, 29, 36, 37, 39, 47, 51] for extensions in the $L^{p}$ Brunn–Minkowski theory and for more general sets. Recently, Milman and Yehudayoff [42] established isoperimetric inequalities for affine quermassintegrals. These new inequalities are significantly stronger than the isoperimetric inequalities for the classic quermassintegrals, thereby settling a long-standing conjecture of Lutwak [35].
@@
More recently, Haddad and Ludwig [25] established the following sharp affine fractional Sobolev inequality, which is significantly stronger than the classical inequality (1.4).
Affine fractional Sobolev inequalities: For $1<\alpha<0$ and $f\in W^{-\alpha,1}(\mathbb{R}^{n})$, that is, $f$ with finite $\|f\|_{W^{-\alpha,1}(\mathbb{R}^{n})}$,
$\begin{split}2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha}\Big\|\frac{f(\cdot+tu)-f(\cdot)}{t}\Big\|_{1}dt\Big)^{\frac{n}{\alpha}}\Big)^{\frac{\alpha}{n}}\\ \leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy.\end{split}$ (1.7)
There is equality in the first inequality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid. There is equality in the second inequality for radially symmetric functions.
In the limit as $\alpha\rightarrow-1^{+}$, the inequality (1.7) recovers the well-known affine Sobolev inequality established by Zhang [51], which is an affine version of the classic $L^{1}$ Sobolev inequality. For a Sobolev function $f$, Zhang [51] defined the polar projection body $\Pi^{*}f$ by
$\|\xi\|_{\Pi^{*}f}=\frac{1}{2}\int_{\mathbb{R}^{n}}|\nabla f(x)\cdot\xi|dx,~~\xi\in\mathbb{S}^{n-1},$
where $\|\cdot\|_{K}$ denotes the gauge function of $K$. Wang [47] extended the definition to functions of bounded variation and obtained the generalized Sobolev-Zhang inequality. The extended version corresponds to the functional form of the Petty projection inequality. Using (1.7), Haddad and Ludwig derived the fractional Petty projection inequalities by taking $f=\operatorname{1}_{K}$.
Motivated by the same central ideas as in their earlier work [25], Haddad and Ludwig [26, 27] established the affine fractional $L^{p}$ Sobolev inequalities as well as affine HLS inequalities. For further results on affine functional inequalities, we refer the reader to [9, 11, 22].
The following Theorem 2 presents an affine version of our functional isoperimetric inequality for chord power integrals, and thus strengthens Theorem 1.
@@
For $\alpha\in(0,n)$ and non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt\Big)^{\frac{n}{\alpha}}\Big)^{\frac{\alpha}{n}}$ (1.8)
@@
In Section 4, we introduce an anisotropic version of the right side of (1.8), which gives rise to a new star-shaped set ${\rm L}_{\alpha}f$, defined via its radial function by
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt,~~u\in\mathbb{S}^{n-1}.$
With the definition, the first inequality in Theorem 2 can be written as
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$ (1.9)
Since both sides of (1.9) are invariant under translations of $f$ and
@@
for every volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, it follows that inequality (1.9) is indeed affine invariant.
According to (1.1) and (1.2) the isoperimetic inequality for $I_{\alpha+1}(K)$ reverses when $\alpha>n$. Theorem 1 establishes the functional analogue for $\alpha\in(0,n)$. However, the case $\alpha>n$ does not admit a straightforward reversal, which will be discussed in detail in Section 6.
The next result presents the affine inequality we obtain for $\alpha>n$, where ${\rm L}_{\alpha}f$ has been introduced above and $L^{\infty}(\mathbb{R}^{n})$ is the space of essentially bounded measurable functions, equipped with the essential supremum norm $\|\cdot\|_{\infty}$. By eliminating the intermediate term, it serves as the functional analogue of (1.1) for $\alpha>n$.
Theorem 3.
Let $\alpha>n$ and $f$ be a non-zero, non-negative function in $L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$. Then
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
In Section 7 and Section 8, we relate the set ${\rm L}_{\alpha}f$ to radial mean bodies originally introduced by Gardner and Zhang [18]. For $\alpha>-1$ and a convex body $K\subset\mathbb{R}^{n}$, the radial $\alpha$-th mean body ${\rm R}_{\alpha}K$ is defined via its radial function by
$\rho_{{\rm R}_{\alpha}K}(u)=\Big(\frac{1}{|K|}\int_{K}\rho_{K-x}(u)^{\alpha}dx\Big)^{1/\alpha},\quad\quad\alpha\neq 0,$
@@
$\log\rho_{{\rm R}_{0}K}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx,\quad\quad\alpha=0.$
For $\alpha\in(-1,0)$, Haddad and Ludwig [25] showed that the radial mean body ${\rm R}_{\alpha}K$ is proportional to $\Pi^{*,-\alpha}\operatorname{1}_{K}$, where $\Pi^{*,-\alpha}\operatorname{1}_{K}$ denotes the fractional polar projection body arsing from the affine fractional Sobolev inequality (1.7) (see Section 2.6 for details). Of particular interest, for $f$ with bounded variation, Haddad and Ludwig [25] proved that
$\lim_{\alpha\rightarrow-1^{+}}(1+\alpha)\|u\|_{\Pi^{*,-\alpha}f}^{-\alpha}=\|u\|_{\Pi^{*}f}.$
In Section 7, we show that ${\rm R}_{\alpha}K$ is proportional to ${\rm L}_{\alpha}\operatorname{1}_{K}$ when $\alpha>n$. Motivated by this relation, we introduce an explicit definition of the functional counterpart ${\rm R}_{\alpha}f$, based on the fractional polar projection body $\Pi^{*,-\alpha}f$ and the star-shaped set ${\rm L}_{\alpha}f$. We note that Haddad and Ludwig [27] also used this notation for radial mean bodies, which in their setting are defined via the fractional $L^{2}$ polar projection body and the star-shaped sets appearing in affine HLS inequalities. Here, we use the same notation ${\rm R}_{\alpha}f$ for an $L^{1}$- extension to keep the notation simple, even though the definition in [27] is slightly different.
In our setting, in particular, for a log-concave function $f\in L^{1}(\mathbb{R}^{n})$, we relate ${\rm R}_{\alpha}f$ to the radial mean bodies of its superlevel sets $\{f\geq r\}$. More precisely, for $\alpha>-1$,
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r),\quad\quad\quad\alpha\neq 0,$
and
$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)d\mu_{f}(r),\quad\quad\alpha=0,$
where $d\mu_{f}(r)=\frac{|\{f\geq r\}|}{\|f\|_{1}}dr$.
This unexpected relation naturally leads to the definition of ${\rm R}_{\infty}f$,
${\rm R}_{\infty}f={\rm R}_{\infty}~{\rm supp}(f)={\rm D}~{\rm supp}(f),$
where $f\in L^{1}(\mathbb{R}^{n})$ is a log-concave function with compact support, denoted by ${\rm supp}(f)$. The notation ${\rm D}K$ is the difference body of $K$, defined by ${\rm D}K=\{x-y:x,y\in K\}$. This observation yields the following monotonicity property:
@@
for $1<\alpha<\beta<\infty$. When $f=\operatorname{1}_{K}$, these relations reduce to the geometric case of Gardner and Zhang [18].
Gardner and Zhang also established the reversed inclusion in [18, Theorem 5.5], which requires a more involved argument. For a convex body $K\subset\mathbb{R}^{n}$, they proved
@@
for $1<\alpha<\beta<\infty$. In each inclusion, equality holds if and only if $K$ is a simplex, denoted by $\triangle$. Here $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$, where $B(\alpha+1,n)$ denotes the Beta function.
In (1.10), when $\beta=n$, the left side inclusion implies the Rogers-Shephard inequality [44], which is
$|{\rm D}K|\leq\binom{2n}{n}|K|.$
When $\alpha=n$, the right side inclusion yields Zhang’s inequality [50] (see also Ball [5]):
which is a reversed form of Petty projection inequality.
To extend these geometric inequalities to their functional analogues, there are two natural ways to associate a convex body with a log-concave function:

### v2

# Affine chord Sobolev inequalities and radial mean bodies for functions

@@ Abstract.
Affine isoperimetric inequalities for the functional radial mean bodies are derived from the new affine chord Sobolev inequalities, which extend the recent affine isoperimetric inequalities of Haddad and Ludwig [24] from convex bodies to functions. The affine chord Sobolev inequalities further imply a strengthening of the Euclidean chord Sobolev inequalities introduced in [4]. Moreover, for $s$-concave functions $f$ with compact support and $s>0$, a parameter-dependent monotonicity property of the functional radial mean body ${\rm R}_{\alpha}f$ is obtained: ${\rm R}_{\alpha}f\subset{\rm R}_{\beta}f$ for $1<\alpha<\beta$, and, after suitable normalization, the reverse inclusion also holds. These sharp results generalize the corresponding monotonicity for geometric radial mean bodies established by Gardner and Zhang [18].
@@ 1. Introduction
Affine inequalities, invariant under translations and volume-preserving linear transformations, often strengthen classical Euclidean inequalities. A seminal example is Zhang’s affine Sobolev inequality [48], which significantly improves the classical Sobolev inequality. Wang [44] later extended this inequality to functions of bounded variation, yielding a generalized Petty projection inequality (see also [48]), itself a strengthening of the classical isoperimetric inequality. Recently, Milman and Yehudayoff [39] established affine isoperimetric inequalities for affine quermassintegrals, settling a long-standing conjecture of Lutwak [33]. For further developments on affine inequalities, see, for example, [7, 10, 12, 18, 20, 21, 22, 23, 34, 35, 36].
@@
More recently, Haddad and Ludwig [24, 25, 26] developed a new approach to study affine inequalities. They proved the following sharp affine fractional Sobolev inequalities in [24], which are stronger than their Euclidean counterparts by Almgren and Lieb [2] (see also [15]).
Affine fractional Sobolev inequalities: For $\alpha\in(-1,0)$ and $f\in W^{-\alpha,1}(\mathbb{R}^{n})$,
$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}$ $r^{\alpha-1}\|f(\cdot)-f(\cdot+ru)\|_{1}dr\Big)^{\frac{n}{\alpha}}du\Big)^{\frac{\alpha}{n}}$ (1.1)
$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,$
where $\mathbb{S}^{n-1}$ is the unit sphere in $\mathbb{R}^{n}$ and $\omega_{n}$ denotes the volume of the unit ball in $\mathbb{R}^{n}$. The sharp constant $\sigma_{n,\alpha}$ is given by
$\sigma_{n,\alpha}=\frac{n2^{\alpha}\pi^{\frac{n-\alpha-1}{2}}\Gamma(\frac{\alpha+1}{2})\Gamma(\frac{n}{2}+1)^{\frac{\alpha}{n}}}{|\alpha|\Gamma(\frac{n+\alpha+1}{2})}.$ (1.2)
Here $\Gamma$ denotes the gamma function, $\|f\|_{p}:=(\int_{\mathbb{R}^{n}}f(x)^{p}dx)^{1/p}$ denotes the $L^{p}$ norm of $f$ and $W^{-\alpha,1}(\mathbb{R}^{n})$ is the space of integrable functions $f$ such that the right-most side of (1.1) is finite.
Eliminating the intermediate term in (1.1) gives the fractional Sobolev inequalities, which converge to the classical Sobolev inequality when $\alpha\rightarrow-1^{+}$, as shown by Bourgain, Brezis and Mironescu [8]. Haddad and Ludwig [24] showed that (1.1) becomes the affine Sobolev inequality in the limit $\alpha\rightarrow-1^{+}$. The fractional Sobolev inequalities have been extensively studied over the past decades (see, e.g., [1, 8, 13, 27, 30, 31]). In very recent work by the authors [4], the following sharp chord Sobolev inequalities for positive $\alpha$ are established, where, for $E\subset\mathbb{R}^{n}$, the characteristic function $\operatorname{1}_{E}(x)$ takes value $1$ with $x\in E$ and 0 otherwise.
Chord Sobolev inequalities for $\boldsymbol{\alpha>0}$: For $\alpha\in(0,n)$ and non-negative $f\in L^{\frac{n}{n+\alpha}}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$ (1.3)
For $\alpha>n$ and non-zero, non-negative $f\in L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy,$ (1.4)
where $\sigma_{n,\alpha}$ is given by (1.2). In both cases, equality holds if and only if $f$ is a constant multiple of the characteristic function of a ball.
In [4], fractional Sobolev inequalities, together with (1.3) and (1.4), form a complete series called the chord Sobolev inequalities. When $f=\operatorname{1}_{K}$ is the characteristic function of a convex body $K\subset\mathbb{R}^{n}$ (a convex compact set with non-empty interior), these inequalities reduce to the isoperimetric inequalities for chord power integrals; see [4] for details. Chord power integrals have been studied extensively in recent years [4, 11, 19, 37, 45, 46].
One of the main aims of this paper is to prove the following affine chord Sobolev inequalities for $\alpha>0$, which strengthen their Euclidean counterparts (1.3) and (1.4). Here $f$ is radially symmetric if its value depends only on $|x|$.
@@
For $\alpha\in(0,n)$ and non-negative $f\in L^{\frac{n}{n+\alpha}}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}r^{\alpha-1}$ $\big\|\min\{f(\cdot),f(\cdot+ru)\}\big\|_{1}dr\Big)^{\frac{n}{\alpha}}du\Big)^{\frac{\alpha}{n}}$ (1.5)
@@
Theorem 2.
For $\alpha>n$ and non-zero, non-negative $f\in L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}r^{\alpha-1}$ $\big\|\min\{f(\cdot),f(\cdot+ru)\}\big\|_{1}dr\Big)^{\frac{n}{\alpha}}du\Big)^{\frac{\alpha}{n}}$ (1.6)
$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
The affine fractional Sobolev inequalities (1.1), together with (1.5) and (1.6), are called affine chord Sobolev inequalities. To prove Theorem 1 and Theorem 2, we introduce an anisotropic version of the right-most side of (1.5) (see, e.g., [24, 30, 31]), which gives rise to a new star-shaped set ${\rm L}_{\alpha}f$, defined via its radial function by
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr,~~u\in\mathbb{S}^{n-1},$
(see Section 3 for details). The first inequality in Theorem 1 can then be written as
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}},$ (1.7)
where $|{\rm L}_{\alpha}f|$ denotes the volume of ${\rm L}_{\alpha}f$. Since both sides of (1.7) are invariant under translations of $f$ and
@@
for every volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, it follows that inequality (1.7) is affine invariant.
For a convex body $K\subset\mathbb{R}^{n}$, the star-shaped set ${\rm L}_{\alpha}\operatorname{1}_{K}$ is proportional to the radial mean body ${\rm R}_{\alpha}K$ of Gardner and Zhang [18], as will be discussed in Section 5. Haddad and Ludwig [24] proved affine isoperimetric inequalities for ${\rm R}_{\alpha}K$ with $\alpha>-1$, stating that
$\frac{|{\rm R}_{\alpha}K|}{|K|}$ $\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha\in(-1,n),$ (1.8)
$\frac{|{\rm R}_{\alpha}K|}{|K|}$ $\geq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha>n.$
Equality holds if and only if $K$ is an ellipsoid.
In [4], the authors extended the notion ${\rm R}_{\alpha}K$ to log-concave functions $f\in L^{1}(\mathbb{R}^{n})$ by
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq t\}}(u)^{\alpha}d\mu_{f}(t),~\alpha\neq 0,$ (1.9)
@@
$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq t\}}(u)d\mu_{f}(t),$
where $d\mu_{f}(t)=\frac{|\{f\geq t\}|}{\|f\|_{1}}dt$ is a probability measure on $\mathbb{R}$. We note that Haddad and Ludwig [26] also use the notation ${\rm R}_{\alpha}f$ for the $\alpha$-th radial mean body, but with a different definition. The present definition can be viewed as an $L^{1}$ extension to their definitions.
It was shown in [4] that ${\rm R}_{\alpha}f$ is proportional to ${\rm L}_{\alpha}f$ for $\alpha>0$, while for $\alpha\in(-1,0)$ it is proportional to the fractional polar projection body $\Pi^{*,-\alpha}f$ arising from the affine fractional Sobolev inequalities (1.1). Consequently, the affine chord Sobolev inequalities yield the following affine isoperimetric inequalities for ${\rm R}_{\alpha}f$: for a log-concave function $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$,
$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}$ $\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha\in(-1,n),$ (1.10)
$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{1}\|f\|_{\infty}^{-1}}$ $\geq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha>n,$
with equality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid. In particular, when $f=\operatorname{1}_{K}$ is the characteristic function of a convex body, (1.10) reduces to the geometric inequality (1.8).
Of particular interest, for log-concave $f\in L^{1}(\mathbb{R}^{n})$, Haddad and Ludwig [24] proved
$\lim_{\alpha\rightarrow-1^{+}}(1+\alpha)\|f\|_{1}\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\rho_{\Pi^{*}f}(u)^{-1},$
where $\Pi^{*}f$ is the polar projection body of $f$ introduced by Zhang [48] and extended by Wang [44] (see Section 2.5). For the other endpoint $\alpha\to\infty$, the monotonicity of $L^{p}$ norms with respect to $p$, together with (1.9), leads to the definition of ${\rm R}_{\infty}f$:
${\rm R}_{\infty}f={\rm R}_{\infty}\,{\rm supp}(f)={\rm D}\,{\rm supp}(f),$
where ${\rm D}K:=\{x-y:x,y\in K\}$ denotes the difference body, and we set ${\rm R}_{\infty}f=\mathbb{R}^{n}$ if ${\rm supp}(f)=\mathbb{R}^{n}$.
In Section 5, we prove the following monotonicity for log-concave $f\in L^{1}(\mathbb{R}^{n})$:
@@
$\|f\|_{\infty}\Pi^{*}f\subset\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$ (1.11)
for $1<\alpha<\beta<\infty$. When $f=\operatorname{1}_{K}$, these reduce to the geometric case of Gardner and Zhang [18].
Gardner and Zhang [18] also established the reverse inclusion for radial mean bodies, which is more involved. For a convex body $K\subset\mathbb{R}^{n}$, they proved
@@
for $1<\alpha<\beta<\infty$, where $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$ and $B(\cdot,\cdot)$ is the Beta function. Equality holds in each inclusion if and only if $K$ is a simplex. When $\beta=n$, the left-most inclusion gives the Rogers–Shephard inequality [41], while for $\alpha=n$, the right-most inclusion yields Zhang’s inequality [47] (see also Ball [5]).
Building on the functional radial mean body ${\rm R}_{\alpha}f$ and (1.11), we aim to extend the reverse inclusion (1.12) from sets to functions. There are two natural ways to associate a convex body with a log-concave function: